# TrackGo: A Flexible and Efficient Method for Controllable Video Generation

### v1

@@ Abstract
Recent years have seen substantial progress in diffusion-based controllable video generation. However, achieving precise control in complex scenarios, including fine-grained object parts, sophisticated motion trajectories, and coherent background movement, remains a challenge. In this paper, we introduce TrackGo, a novel approach that leverages free-form masks and arrows for conditional video generation. This method offers users with a flexible and precise mechanism for manipulating video content. We also propose the TrackAdapter for control implementation, an efficient and lightweight adapter designed to be seamlessly integrated into the temporal self-attention layers of a pretrained video generation model. This design leverages our observation that the attention map of these layers can accurately activate regions corresponding to motion in videos. Our experimental results demonstrate that our new approach, enhanced by the TrackAdapter, achieves state-of-the-art performance on key metrics such as FVD, FID, and ObjMC scores. The project page of TrackGo can be found at: https://zhtjtcz.github.io/TrackGo-Page/
@@
Introduction
With the rapid development of diffusion models (Ho, Jain, and Abbeel 2020; Song, Meng, and Ermon 2020; Nichol and Dhariwal 2021; Dhariwal and Nichol 2021; Song et al. 2020), video generation has witnessed significant progress, with the quality of generated videos continuously improving. Unlike text-based (Blattmann et al. 2023b; Zhang et al. 2023; Guo et al. 2023b) or image-based (Blattmann et al. 2023a) video generation, controllable video generation (Hu et al. 2023; Yin et al. 2023; Wu et al. 2024) focuses on achieving precise control over object movement and scene transformations in generated videos. This capability is particularly valuable in industry such as film production and cartoon creation.
@@ Introduction
We introduce TrackGo, a novel framework for controllable video generation that leverages user inputs to direct the generation of video sequences. TrackGo uses free-form masks and arrows provided by users to define target regions and movement trajectories, respectively. This approach consists of two stages: Point Trajectories Generation and Conditional Video Generation. In the first stage, TrackGo automatically extracts point trajectories from the user-defined masks and arrows. These trajectories serve as precise blueprints for video generation. In the second stage, we use the Stable Video Diffusion Model (SVD) (Blattmann et al. 2023a) as the base model, accompanied by an encoder that encodes the motion information. To ensure that the guidance is precisely interpreted by our model, we introduce the novel TrackAdapter. This adapter effectively modifies the existing temporal self-attention layers of a pre-trained video generation model to accommodate new conditions, enhancing the model’s control over the generated video.
@@
Diffusion models (Ho, Jain, and Abbeel 2020) have made great progress in the field of text-to-image generation (Betker et al. 2023; Peebles and Xie 2023; Rombach et al. 2022; Saharia et al. 2022; Zhang, Rao, and Agrawala 2023; Xing et al. 2023), which directly promotes the progress of basic video diffusion, and many excellent works (Blattmann et al. 2023a, b; Chen et al. 2024; Guo et al. 2023b; Henschel et al. 2024; Ho et al. 2022; Zhang et al. 2023; Wang et al. 2023a; Zeng et al. 2023) have emerged. Early works such as VLDM (Blattmann et al. 2023b) and AnimateDiff (Guo et al. 2023b) try to insert temporal layers to complete the generation of text to video. Recent models benefit from the stability of diffusion-based trained models. I2VGen-XL (Zhang et al. 2023), Stable Video Diffusion (SVD) (Blattmann et al. 2023a) achieve surprising results on text-to-video generation and image-to-video, respectively, with large-scale high-quality data. While these models are capable of producing high-quality videos, they primarily depend on coarse-grained semantic guidance from text or image prompts, which can lead to actions that do not align with the user’s intentions.
@@
To broaden the scope of controllable video synthesis  (Wang et al. 2024b; Guo et al. 2023a; Ma et al. 2024; Wang et al. 2023c; Yin et al. 2023; Wang et al. 2024a; Wu et al. 2024; Tu et al. 2023; Jiang et al. 2023; Qi et al. 2023) and enhance its general applicability, several approaches incorporate control signals into pre-trained video diffusion models. The work of AnimateDiff (Guo et al. 2023b) endeavors, to integrate LoRA (Low-Rank Adaptation) (Hu et al. 2021) within temporal attention mechanisms to grasp camera motion learning. Nevertheless, LoRA’s potential shortcomings emerge when tasked with thoroughly manipulating and synthesizing sophisticated movements, as its learning capacity and predictive prowess might be constrained by repetitive motion patterns present in the training dataset. Innovations like MotionCtrl (Wang et al. 2023c) and DragNUWA (Yin et al. 2023) encode sparse optical flow into dense optical flow as guidance information to inject into the diffusion model to control object motion. Boximator (Wang et al. 2024a), on the other hand, pioneers motion regulation by correlating objects with bounding boxes, capitalizing on the model’s innate tracking capabilities. Fundamentally, this approach employs dual trajectory sets defined by the bounding box’s upper-left and lower-right vertices. A caveat arises from the boxes occasionally misinterpreting background details as part of the foreground, which can inadvertently taint output accuracy. DragAnything (Wu et al. 2024) has introduced an approach that employs masks to pinpoint a central point, subsequently generating a Gaussian map for tracking this center to produce a guiding trajectory for model synthesis. It’s crucial to note, however, that not all masked areas denote regions of motion. And relying on the ControlNet (Zhang, Rao, and Agrawala 2023) structure makes it difficult to achieve satisfactory results in terms of efficiency.
@@
Our task is motion-controllable video generation. For an input image $\bm{I}\in\mathbb{R}^{H\times W\times 3}$, and point trajectories $\bm{P}\in\mathbb{R}^{L\times H\times W\times 3}$ extracted from arrows describes the trajectories information, generating a video $\bm{V}\in\mathbb{R}^{L\times H\times W\times 3}$ in line with the trajectories, where $L$ denotes the length of video.
@@
In inference, when the user provides the first frame picture, the masks of the editing area, and the corresponding arrows. Our approach can convert the user’s input masks and arrows into point trajectories $\bm{P}$ through preprocessing, as shown in Fig. 3.
@@
where $\bm{Q}^{{}^{\prime}}=f\bm{W}_{q}^{{}^{\prime}},\bm{K}^{{}^{\prime}}=f\bm{W}_{q}^{{}^{\prime}}$ are the query, key matrices of the TrackAdapter, $f$ is the compressed representation of $\bm{P}$, obtained from Eq. 1.
@@
Finally, we get the temporal self-attention output of the current block: $\bm{O}=(\bm{A}+\bm{A}^{{}^{\prime}})\bm{V}=\bm{A}^{{}^{\prime\prime}}\bm{V}$, where $\bm{Q}=\bm{X}\bm{W}_{q},\bm{K}=\bm{X}\bm{W}_{q},\bm{V}=\bm{X}\bm{W}_{v}$ are the query, key, and value matrices of the temporal self-attention operation, $\bm{X}$ is input feature and $\bm{A^{\prime\prime}}$ is final attention map. We complete the separation of the corresponding area and the unspecified area during the calculation of attention.
@@
In order to adapt the original temporal self-attention to the new input mode quickly and to accelerate the model’s convergence, we design an attention map based loss function. We gather attention maps from different blocks to get a set $\mathcal{C}$, which contains 16 different attention maps. For $\bm{A}_{q}\in\mathcal{C}$, $\bm{A}_{q}(x,y)$ denotes the temporal self-attention map between frame $x$ and frame $y$ at block $q$. The purpose of the attention loss is to suppress the area corresponding to the mask in the original branch’s attention map, i.e. the motion area,
@@
During inference, we set the intensity of the unspecified area to $\tau$, that is, we set the part of the Eq  3 less than $\alpha$ to $\tau$. Users can adjust $\tau$ to control the movement of the unspecified area in cases where it needs to move synchronously with the foreground movement or when sensory interference from the unspecified area needs to be mitigated. This feature will greatly enhance the creation of fluid and highly synchronized motion videos.
@@
Table 1: Qualititative comparisons of our approach with DragAnything and DragNUWA. All three methods are based on the same basic model called SVD. Bolded values indicate the best scores in each column.
@@
Our test set comprised the VIPSeg validation set along with an additional 300-video subset from our internal validation dataset. Notably, all videos in the VIPSeg dataset are formatted in a 16:9 aspect ratio. To maintain consistency, we adjusted the resolution of all videos in the validation sets to $1024\times 576$, unlike $256\times 256$ used in DragAnything. For evaluation purposes, we extracted the trajectories from the first 14 frames of each video in the test set.
@@
We measure video quality using FVD (Unterthiner et al. 2018) (Fréchet Video Distance) and image quality using FID (Heusel et al. 2017). We compare our methods with DragNUWA and DragAnything, which can also use trajectory information as conditional input. Following DragAnything, ObjMC is used to evaluate the motion control performance by computing the Euclidean distance between the predicted and ground truth trajectories.
@@
Quantitative comparisons of our method with baseline methods are shown in Table 1. We test all models on VIPSeg validation set and internal validation set. From the results, we can see that TrackGo outperforms all other approaches across all metrics, indicating that our approach can produce videos with higher visual quality and is more faithful to the input motion control. We also compared the model parameters and inference speed of the three approaches. Since all three approaches use the same base model, our comparison focused exclusively on the total weights of the modules newly added to the base model. To assess the model’s inference speed, we conducted 100 inference tests for each approach using identical input data on an NVIDIA A100 GPU. The results demonstrate that our approach not only delivers the best visual quality but also achieves the fastest inference speed, all while requiring the fewest additional parameters.
@@
We present a visualization comparison with DragAnything and DragNUWA, as shown in Fig. 4. We can have the following observations: First, DragNUWA struggles with perceiving the control area, which may lead to incomplete or inaccurate optical flow. In case (b), the planet is not correctly perceived, and in case (a), the movement of the gun is also incorrect. In case (c), although the optical flow of the train is successfully predicted, the absence of optical flow in the smoke results in a jarring visual effect. Second, DragAnything also faces difficulties with the movement of partial or fine-grained objects. As illustrated in case (a), only the gun and Mario’s hand should move; however, Mario’s entire position also shifts unexpectedly. A similar problem occurs in case (b). Moreover, DragAnything struggles to produce a harmonious background. In case (c), the smoke does not follow the moving train. In contrast, our proposed TrackGo can generate videos where the movement of the target region is precisely aligned with the user input, while maintaining the consistency and harmony of the background. This capability significantly enhances the visual quality and coherence of the generated videos, demonstrating the effectiveness of TrackGo. More cases of our approach can be found in Fig. 1.
@@
Our model possesses the capability to adjust the intensity of motion in unspecified areas through specific parameters, which is illustrated in the Fig. 5. We define the motor inhibition intensity in unspecified areas with the parameter $\tau$. As shown in Fig. 5, when the hands move, it is necessary for the steering wheel to move correspondingly to enhance the realism of the output. When $\tau=-10^{4}$, the motion in unspecified areas is significantly suppressed, allowing only the hands to follow their trajectory while other parts remain static. This often results in a distortion of the overall video. When $\tau=-10^{2}$, the unspecified area is less suppressed, allowing hands and the steering wheel to move simultaneously. However, this setting can still produce disharmonious outcomes, such as the car logo on the steering wheel remaining static. When $\tau=0$, the unspecified areas are allowed to move freely, which typically results in a more cohesive and harmonious video. Nevertheless, not all movements in the unspecified areas are desirable, and excessive movement can damage the video quality. Therefore, it is crucial to carefully manage the suppression level to balance realism with artistic control.
Camera Motion.
Ablation Study
@@
User Study.

### v2

@@ Abstract
Recent years have seen substantial progress in diffusion-based controllable video generation. However, achieving precise control in complex scenarios, including fine-grained object parts, sophisticated motion trajectories, and coherent background movement, remains a challenge. In this paper, we introduce TrackGo, a novel approach that leverages free-form masks and arrows for conditional video generation. This method offers users with a flexible and precise mechanism for manipulating video content. We also propose the TrackAdapter for control implementation, an efficient and lightweight adapter designed to be seamlessly integrated into the temporal self-attention layers of a pretrained video generation model. This design leverages our observation that the attention map of these layers can accurately activate regions corresponding to motion in videos. Our experimental results demonstrate that our new approach, enhanced by the TrackAdapter, achieves state-of-the-art performance on key metrics such as FVD, FID, and ObjMC scores.
Introduction
With the rapid development of diffusion models (Ho, Jain, and Abbeel 2020; Song, Meng, and Ermon 2020; Dhariwal and Nichol 2021; Song et al. 2020), video generation has witnessed significant progress, with the quality of generated videos continuously improving. Unlike text-based (Blattmann et al. 2023b; Zhang et al. 2023; Guo et al. 2023b) or image-based (Blattmann et al. 2023a) video generation, controllable video generation (Hu et al. 2023; Yin et al. 2023; Wu et al. 2024) focuses on achieving precise control over object movement and scene transformations in generated videos. This capability is particularly valuable in industry such as film production and cartoon creation.
@@ Introduction
We introduce TrackGo, a novel framework for controllable video generation that leverages user inputs to direct the generation of video sequences. TrackGo uses free-form masks and arrows provided by users to define target regions and movement trajectories, respectively, as illustrated in Fig. 2. This approach consists of two stages: Point Trajectories Generation and Conditional Video Generation. In the first stage, TrackGo automatically extracts point trajectories from the user-defined masks and arrows. These trajectories serve as precise blueprints for video generation. In the second stage, we use the Stable Video Diffusion Model (SVD) (Blattmann et al. 2023a) as the base model, accompanied by an encoder that encodes the motion information. To ensure that the guidance is precisely interpreted by our model, we introduce the novel TrackAdapter. This adapter effectively modifies the existing temporal self-attention layers of a pre-trained video generation model to accommodate new conditions, enhancing the model’s control over the generated video.
@@
Diffusion models (Ho, Jain, and Abbeel 2020) have made great progress in the field of text-to-image generation (Betker et al. 2023; Peebles and Xie 2023; Rombach et al. 2022; Saharia et al. 2022; Zhang, Rao, and Agrawala 2023; Xing et al. 2023), which directly promotes the progress of basic video diffusion, and many excellent works (Chen et al. 2024; Henschel et al. 2024; Ho et al. 2022; Zhang et al. 2023; Wang et al. 2023a; Zeng et al. 2023) have emerged. Early works such as VLDM (Blattmann et al. 2023b) and AnimateDiff (Guo et al. 2023b) try to insert temporal layers to complete the generation of text to video. Recent models benefit from the stability of diffusion-based trained models. I2VGen-XL (Zhang et al. 2023), Stable Video Diffusion (SVD) (Blattmann et al. 2023a) achieve surprising results on text-to-video generation and image-to-video, respectively, with large-scale high-quality data. While these models are capable of producing high-quality videos, they primarily depend on coarse-grained semantic guidance from text or image prompts, which can lead to actions that do not align with the user’s intentions.
@@
To broaden the scope of controllable video synthesis  (Wang et al. 2024b; Guo et al. 2023a; Ma et al. 2024; Wang et al. 2023c; Yin et al. 2023; Wang et al. 2024a; Wu et al. 2024; Tu et al. 2023; Jiang et al. 2023; Qi et al. 2023) and enhance its general applicability, several approaches incorporate control signals into pre-trained video diffusion models. The work of AnimateDiff endeavors, to integrate LoRA (Low-Rank Adaptation) (Hu et al. 2021) within temporal attention mechanisms to grasp camera motion learning. Nevertheless, LoRA’s potential shortcomings emerge when tasked with thoroughly manipulating and synthesizing sophisticated movements, as its learning capacity and predictive prowess might be constrained by repetitive motion patterns present in the training dataset. Innovations like MotionCtrl (Wang et al. 2023c) and DragNUWA (Yin et al. 2023) encode sparse optical flow into dense optical flow as guidance information to inject into the diffusion model to control object motion. Boximator (Wang et al. 2024a), on the other hand, pioneers motion regulation by correlating objects with bounding boxes, capitalizing on the model’s innate tracking capabilities. Fundamentally, this approach employs dual trajectory sets defined by the bounding box’s upper-left and lower-right vertices. A caveat arises from the boxes occasionally misinterpreting background details as part of the foreground, which can inadvertently taint output accuracy. DragAnything (Wu et al. 2024) has introduced an approach that employs masks to pinpoint a central point, subsequently generating a Gaussian map for tracking this center to produce a guiding trajectory for model synthesis. It’s crucial to note, however, that not all masked areas denote regions of motion. And relying on the ControlNet (Zhang, Rao, and Agrawala 2023) structure makes it difficult to achieve satisfactory results in terms of efficiency.
@@
Our task is motion-controllable video generation. For an input image $\bm{I}\in\mathbb{R}^{H\times W\times 3}$, and point trajectories $\bm{P}\in\mathbb{R}^{L\times H\times W\times 3}$ extracted from arrows describes the trajectories information, generating a video $\bm{V}\in\mathbb{R}^{L\times H\times W\times 3}$ in line with the trajectories, where $L$ is the length of video.
@@
In inference, when the user provides the first frame picture, the masks of editing area, and corresponding arrows. Our approach can convert the masks and arrows into point trajectories $\bm{P}$ through preprocessing, as shown in Fig. 3.
@@
where $\bm{Q}^{{}^{\prime}}=f\bm{W}_{q}^{{}^{\prime}},\bm{K}^{{}^{\prime}}=f\bm{W}_{k}^{{}^{\prime}}$ are the query, key matrices of the TrackAdapter, $f$ is the compressed representation of $\bm{P}$, obtained from Eq. 1.
@@
Finally, we get the temporal self-attention output of the current block: $\bm{O}=(\bm{A}+\bm{A}^{{}^{\prime}})\bm{V}=\bm{A}^{{}^{\prime\prime}}\bm{V}$, where $\bm{Q}=\bm{X}\bm{W}_{q},\bm{K}=\bm{X}\bm{W}_{k},\bm{V}=\bm{X}\bm{W}_{v}$ are the query, key, and value matrices of the temporal self-attention operation, $\bm{X}$ is input feature and $\bm{A^{\prime\prime}}$ is final attention map. We complete the separation of the corresponding area and the unspecified area during the calculation of attention.
@@
In order to adapt the original temporal self-attention to the new input mode quickly and to accelerate the model’s convergence, we design an attention map based loss function. We gather attention maps from different temporal self-attention layers to get a set $\mathcal{C}$, which contains 16 different attention maps. For $\bm{A}_{q}\in\mathcal{C}$, $\bm{A}_{q}(x,y)$ denotes the temporal self-attention map between frame $x$ and frame $y$ at block $q$. The purpose of the attention loss is to suppress the area corresponding to the mask in the original branch’s attention map, i.e. the motion area,
@@
During inference, we set the intensity of the unspecified area to $\tau$, that is, we set the part of the Eq  11 less than $\alpha$ to $\tau$. Users can adjust $\tau$ to control the movement of the unspecified area in cases where it needs to move synchronously with the foreground movement or when sensory interference from the unspecified area needs to be mitigated. This feature will greatly enhance the creation of fluid and highly synchronized motion videos.
@@
Table 1: Quantitative comparisons of our approach with DragAnything and DragNUWA. All three methods are based on the same basic model called SVD. Bolded values indicate the best scores in each column.
@@
Our test set comprised the VIPSeg validation set along with an additional 300-video subset from our internal validation dataset. Notably, all videos in the VIPSeg dataset are formatted in a 16:9 aspect ratio. To maintain consistency, we adjusted the resolution of all videos in the validation sets to $1024\times 576$. For evaluation purposes, we extracted the trajectories from first 14 frames of each video in the test set.
@@
We measure video quality using FVD (Unterthiner et al. 2018) and image quality using FID (Heusel et al. 2017). We compare our methods with DragNUWA and DragAnything, which can also use trajectory information as conditional input. Following DragAnything, ObjMC is used to evaluate the motion control performance by computing the Euclidean distance between the predicted and ground truth trajectories.
@@
Quantitative comparisons of our method with baseline methods are shown in Table 1. We test all models on VIPSeg validation set and internal validation set. From the results, TrackGo outperforms other approaches across all metrics, producing videos with higher visual quality and better fidelity to input motion control. We also compared the model parameters and inference speed of the three approaches. Since all use the same base model, our comparison focused on the weights of the newly added modules. To assess the model’s inference speed, we conducted 100 inference tests for each approach using identical input data on an NVIDIA A100 GPU. The results demonstrate that our approach not only delivers the best visual quality but also achieves the fastest inference speed, all while requiring the fewest additional parameters.
@@
We present a visualization comparison with DragAnything and DragNUWA, as shown in Fig. 4. We can have the following observations: First, DragNUWA struggles with perceiving the control area, which may lead to incomplete or inaccurate optical flow. In case (b), the planet is not correctly perceived, and in case (a), the movement of the gun is also incorrect. In case (c), while the optical flow of the train is successfully predicted, the absence of flow in the smoke creates a jarring effect. Second, DragAnything struggles with partial or fine-grained object movement. In case (a), only the gun and Mario’s hand should move, but Mario’s entire position shifts unexpectedly, with a similar issue in case (b). DragAnything also fails to produce a harmonious background; in case (c), the smoke doesn’t follow the moving train. In contrast, TrackGo generates videos where target region movement aligns precisely with user input while maintaining background consistency and harmony. This significantly enhances the visual quality and coherence of the videos, showcasing TrackGo’s effectiveness. Additional examples are shown in Fig. 2.
@@
Our model possesses the capability to adjust the intensity of motion in unspecified areas through specific parameters, which is shown in the Fig. 5. We define the motor inhibition intensity in unspecified areas with the parameter $\tau$. As shown in Fig. 5, when the hands move, it is necessary for the steering wheel to move correspondingly to enhance the realism of the output. When $\tau=-10^{4}$, the motion in unspecified areas is significantly suppressed, allowing only the hands to follow their trajectory while other parts remain static. This often results in distortion. When $\tau=-10^{2}$, the unspecified area is less suppressed, allowing hands and the steering wheel to move simultaneously. However, this setting can still produce disharmonious outcomes, such as the car logo on the steering wheel remaining static. When $\tau=0$, the unspecified areas can move freely, which typically results in a more cohesive and harmonious video. Nevertheless, not all movements in the unspecified areas are desirable, and excessive movement can damage the video quality. Therefore, it is crucial to carefully manage the suppression level to balance realism with artistic control.
@@
Conclusion
Acknowledgments
This work was supported by the National Science and Technology Major Project (No. 2022ZD0117800), Young Elite Scientists Sponsorship Program by CAST, and the Fundamental Research Funds for the Central Universities.
Supplementary Material
Overview
This supplementary material is organized into several sections, each providing additional details and analyses related to our work on TrackGo. The sections are structured as follows:
  • In Section A, we present the implementation details of TrackGo, including the default parameters used in the pipeline.
  • In Section B, we provide the experimental settings and details of our internal dataset.
  • In Section C, we conduct a user study to demonstrate the popularity of our method compared to existing approaches.
  • In Section D, we compare our method with the Stable Video Diffusion Model (SVD) (Blattmann et al. 2023a), which serves as the baseline for our method, thereby demonstrating the effectiveness of our proposed TrackAdapter.
  • In Section E, we present additional qualitative results of TrackGo.
  • In Section F, we discuss the limitations of the current version of TrackGo and propose potential improvements.
  • In Section G, we examine the societal impact of TrackGo.
A. Implementation Details of TrackGo
How to Determine $K$. When generating point trajectories from the training dataset, we set the maximum number of significant components, $s$, to 6. The number of points $K$, which ranges from 16 to 64, is determined by the area of the component’s mask, as shown in Eq. 10.
$K=\begin{cases}16&\text{if }p<0.1\\ 16+\frac{(64-16)}{(0.4-0.1)}\cdot(p-0.1)&\text{if }0.1\leq p\leq 0.4\\ 64&\text{if }p>0.4\end{cases}$ (10)
where $p$ represents the proportion of the component’s area to the total image area.
How to Set the Colors of the Control Points. For each control point, we create a visual representation by drawing a square with a 4-pixel side length centered at its coordinates. Colors for these squares are assigned according to a predefined list, correlating with the component each belongs to, and following a specific drawing order.
How to Obtain the Attention Mask. We insert the TrackAdapter into the temporal self-attention of each layer of the 3D U-Net. We obtained the attention map $\bm{A^{\prime}}$ from TrackAdapter and derived a binarized mask through the threshold $\alpha$, as shown in Eq. 11. The $\alpha$ is set to $0.2$.
${\bm{A}_{M}}=\begin{cases}-\inf&\text{if $\bm{A^{{}^{\prime}}}\geq\alpha$ }\\ 0,&\text{if $\bm{A^{{}^{\prime}}}<\alpha$}\end{cases}$ (11)
To prevent overflow during computation with the torch.fp16 data type, we set the value of attention map $\bm{A_{M}}$ to $-10^{4}$ instead of $-inf$.
B. Experimental Settings and Dataset
In this section, we provide additional details about TrackGo’s experimental environment and the internal dataset.
All experiments are performed on a server with 8 NVIDIA A100 GPUs and Intel(R) Xeon(R) Platinum 8336C CPUs. We used the 80GB version of the NVIDIA A100 GPU. During model inference, it takes about 30 seconds to generate a video.
The internal dataset we used to train our model contains approximately 200K video clips, which vary in length from 20 to 600 frames, with a resolution of $1280\times 720$. Most of the video content is real-world. These videos are of high quality and include a variety of motion types, such as vehicle movements, human actions, and various activities. We removed videos from the dataset that were of relatively low quality, had minimal movements, contained discontinuous segments (with transitions or frame jumps), or exhibited significant camera motion. Ultimately, we obtained about 110K video clips as our training dataset.
Figure 6: The results of the user study.
Figure 7: Visualizations of the temporal attention map of SVD and TrackGo. Both approaches use the same image as the input, and the parameters used for inference are the same. With our proposed TrackAdpater, the temporal self-attention layers of TrackGo can better focus on the intended moving object.
C. User Study